# AuxDet: Auxiliary Metadata Matters for Omni-Domain Infrared Small Target Detection

### v1

@@ 3.1 Overall Architecture
As illustrated in Fig. 2, the proposed AuxDet framework establishes a novel multi-modal architecture for omni-domain IRSTD, building upon the Cascade RPN (Vu et al., 2019) as its baseline network. Our architecture introduces three dedicated components that synergistically address cross-domain generalization challenges: (1) An auxiliary metadata processing stream that transforms heterogeneous sensor parameters into latent semantic representations, (2) A hierarchical fusion mechanism for vision-metadata co-adaptation, and (3) A edge-aware refinement pipeline for target detail recovery.
@@
The pipeline initiates with parallel feature extraction pathways, a conventional CNN backbone for visual pattern mining and a metadata pre-encoder that embeds physical sensor characteristics into high-dimensional manifolds through learnable tensor projections. These complementary representations undergo progressive alignment in Multi-Modal Dynamic Modulation Module. The modulated features then enter our Lightweight Edge Enhancement Module, which deploys phased 1D convolutional blocks along spatial-channel dimensions to amplify high-frequency target signatures.
3.2 Auxiliary Feature Processing and Fusion
@@
Given an input image $\mathbf{x}\in\mathbb{R}^{3\times H\times W}$, we extract hierarchical visual features $\{\mathbf{x}_{i}\}_{i=1}^{4}$ through a CNN backbone, where shallow layers ($i=1,2$) preserve critical spatial details for small target localization, and thus become the embedded layers for our core modules. Our multi-modal framework introduces parallel processing of heterogeneous auxiliary metadata =$\mathcal{M}=\{platform,band,resolution\}$ via the Multi-Modal Dynamic Modulation Module, establishing cross-modal correspondences in the latent space.
@@
Heterogeneous Metadata Encoding: Addressing the multi-modal nature of auxiliary information $\mathcal{M}$, we devise a structured encoding strategy. Categorical variables (platform/band) undergo one-hot encoding followed by modality-specific projection:
@@
where $\sigma_{w},\sigma_{h}$ represent dataset-specific normalization factors. The multi-modal embeddings undergo modality-specific projection operations to align with visual feature space:
@@
Cross-Modal Feature Compensation: To address the intrinsic conflict between deep semantic abstraction and fine-grained target preservation, we propose a differential detail amplifier that explicitly models inter-layer geometric discrepancies.
Shallow features $\textbf{x}_{i}$ undergo strided convolution $\mathcal{C}_{\downarrow}$ for spatial alignment, generating pseudo-deep features $\hat{\textbf{x}}_{4}$ that preserve high-frequency gradients:
@@
$\textbf{x}_{e}=\mathcal{F}_{e}(\Delta\textbf{x}),\quad\mathcal{F}_{e}\in\mathbb{R}^{C\times C/4},$ (6)
@@
where $\mathbf{\Psi}_{aux}$ implements three-stage residual MLP blocks. This dual-stream architecture establishes symbiotic relationships between global metadata context and local texture reinforcement, enabling dynamic feature recalibration across omni-domain scenarios.
Our proposed fusion mechanism provides two synergistic benefits: 1) Metadata embeddings z inject domain-specific priors through learnable affine transformations, conditioning network behavior on imaging physics. 2) Compensation features $\textbf{x}_{e}$ preserve high-frequency target signatures via cross-layer gradient propagation, effectively counteracting the texture erosion in deep layers. The combined effect significantly enhances model robustness against domain shifts while maintaining sensitivity to small targets.
3.3 Multi-Modal Dynamic Modulation Module
Our module establishes an instance-adaptive fusion paradigm through conditional feature transformation. Given aligned visual feature $\mathbf{x}_{i}$ and auxiliary feature $\mathbf{a}_{i}$ , we first project them into latent interaction space via average pooling and learnable linear bottleneck $\mathcal{P}_{\theta}\in\mathbb{R}^{C\times D}$ where $D=\lfloor C/2\rfloor$, then progressively refine the representations through two-stage dynamic MLP transformations (Yang et al., 2022). This cascaded architecture enables iterative cross-modal attention:
$\mathbf{y}^{\prime}_{i}=\mathcal{F}_{dyn}^{(2)}\left(\mathcal{F}_{dyn}^{(1)}(\mathcal{P}_{\theta}(\mathbf{y}_{i}))\right)$ (8)
@@ 3.3 Multi-Modal Dynamic Modulation Module
Unlike conventional batch-shared modulation, our per-sample parameterization preserves batch dimension $B$ to enable instance-wise domain adaptation based on auxiliary metadata.
@@
$\mathbf{f}_{sm_{i}}=\sum_{k=1}^{2}\mathcal{G}_{k}\ast\mathbf{f}_{pool_{i}}^{(k)},\quad\mathcal{G}_{k}\in\mathbb{R}^{G\times 1\times 3\times 3}$ (12)
@@
3.4 Lightweight Edge Enhancement Module
A pivotal challenge in IRSTD lies in the vulnerability of small target features to being overwhelmed by intricate background clutter. To counter this, we introduce a Lightweight Edge Enhancement Module aimed at refining fused features and restoring subtle target details, all while minimizing the introduction of additional parameters. This module also mitigates domain sensitivity issues in normalization layers stemming from the diverse nature of infrared imagery.
@@ 3.4 Lightweight Edge Enhancement Module
where ${W}_{3\times 1}$ and ${W}_{1\times 3}$ denote learnable convolutional kernels, $\alpha\in\mathbb{R}$ acts as a gating parameter.
Inspired by Marr-Hildreth edge detection theory, we initialize kernels with discrete Laplacian values to establish stable geometric priors:
@@
This configuration approximates to the 2D Laplacian-of-Gaussian (LoG) filtering:
@@
This formulation ensures intrinsic curvature sensitivity, yielding robust edge responses from the initial training stages. Consequently, our model achieves superior small target awareness, outperforming most baselines (AP50: 73.0%+) within just 6 epochs.
Following the edge convolution, we implement Lightweight Instance Normalization (LW-IN), applied per sample to avoid the domain shift pitfalls of batch normalization, ensuring consistent performance across varied imaging scenarios. The output is expressed as:
@@
where $\mu_{\text{space}}$ and $\sigma_{\text{space}}^{2}$ denote the spatial mean and variance computed across each channel, $\gamma\in\mathbb{R}^{C}$ is the learnable scale parameter, and $\epsilon=10^{-5}$ is used for numerical stability.
4 Experiments
Table 1: Comparison with Other State-of-the-Art Methods on WideIRSTD-Full Dataset
Method WideIRSTD-Full Dataset
Category Venue Backbone FLOPs $\downarrow$ Params. $\downarrow$ AP50 (%) $\uparrow$ Recall (%) $\uparrow$
@@ 4 Experiments
AuxDet (Ours) - ResNet50 0.215T 45.279M 77.9 87.2
4.1 Experimental Settings
1) Datasets: To thoroughly evaluate the effectiveness of the proposed AuxDet framework, we adopt the publicly available WideIRSTD-Full Dataset, a meticulously curated benchmark for IRSTD. This dataset integrates seven prominent public datasets, namely SIRST-V2 (Dai et al., 2023), IRSTD-1K (Zhang et al., 2022b), IRDST (Sun et al., 2023b), NUDT-SIRST (Li et al., 2023), NUDT-SIRST-Sea (Wu et al., 2023), NUDT-MIRSDT (Li et al., 2025), and Anti-UAV (Jiang et al., 2023), augmented by an additional dataset from the National University of Defense Technology. As depicted in Fig. 1, WideIRSTD-Full spans diverse imaging platforms, spectral bands, and target types, encompassing both simulated and real-world data with significant variability, which includes 9,000 training images with corresponding auxiliary information. And we use 7,000 images for training and 2,000 images for validation [^2] https://github.com/YeRen123455/ICPR-Track2-LightWeight/tree/main/dataset/ICPR_Track2/70_20. It stands out due to its extensive auxiliary metadata, including sensor platforms, spectral bands, and resolutions, which are critical for our scene-aware approach. Unlike other datasets, WideIRSTD-Full’s rich contextual information and vast scale make it uniquely suited to assess AuxDet’s generalization across omni-domain scenarios.
2) Implementation Details: The proposed method is implemented using PyTorch 2.0.1 and MMDetection 3.3.0, with an NVIDIA RTX 3090 acceleration via CUDA 11.7. The maximum training duration is 12 epochs, with the learning rate initialized at 0.005 using a linear warmup for the first 500 iterations. The batch size is set as 4 and all images are resized to $1024\times 1024$. Recall and Average Precision (AP50) are employed as primary metrics at 0.5 threshold of IoU.
4.2 Comparison with State-of-the-Arts
We conduct rigorous comparisons with 26 state-of-the-art (SOTA) detectors on the WideIRSTD-Full dataset. The compared methods fall into two categories: general object detection and small target specialized detectors. General object detection methods include one-stage detectors such as RetinaNet (Lin et al., 2017), TOOD (Feng et al., 2021), FCOS (Tian et al., 2019), GFL (Li et al., 2020), ATSS (Zhang et al., 2020b), and AutoAssign (Zhu et al., 2020); two-stage detectors including Faster R-CNN (Ren et al., 2015), Cascade R-CNN (Cai and Vasconcelos, 2019), Grid R-CNN (Lu et al., 2019), Cascade RPN (Vu et al., 2019), Libra R-CNN (Pang et al., 2019), Dynamic R-CNN (Zhang et al., 2020a), and SABL (Wang et al., 2020); as well as end-to-end transformer-based approaches such as Deformable DETR (Zhu et al., 2021), Sparse R-CNN (Sun et al., 2021a), DINO (Zhang et al., 2022a), and DAB-DETR (Liu et al., 2022a). Small target specialized detectors considered include ALCNet (Dai et al., 2021b), NWD (Xu et al., 2022a), RFLA (Xu et al., 2022b), OSCAR (Dai et al., 2023), DNANet (Li et al., 2023), RDIAN (Sun et al., 2023a), DQ-DETR (Huang et al., 2025), EFLNet (Yang et al., 2024), and PConv (Yang et al., 2025).
As shown in Tab. 1, our analysis reveals three principal observations:
@@ 4.2 Comparison with State-of-the-Arts
1) Generic Detector Inadequacy: Despite architectural advances, generic detectors for natural images exhibit catastrophic failures (avg. AP50<45%) due to inherent scale biases. Among these methods, Cascade RPN emerges as the best performer (74.4% AP50) through its multi-stage anchor refinement and adaptive convolution that mitigates feature misalignment. DINO’s contrastive denoising (82.8% Recall) improves small object sensitivity but suffers from transformer’s intrinsic quadratic memory complexity.
@@
2) Limitations of Single-Modal Specialized Models: Domain-specific approaches reveal critical tradeoffs. For example, EFLNet’s adaptive threshold focal loss (73.7% AP50) successfully suppresses background but over-prunes channel dimensions, collapsing on faint targets (68.1% Recall). RFLA’s Gaussian receptive field matching (71.8% AP50) better aligns tiny targets yet requires cascaded attention blocks that bloat parameters (123M vs. ours 45M).
@@
3) AuxDet’s Multi-Modal Breakthrough: Our framework achieves SOTA 77.9% AP50 (+3.5%↑) and 87.2% Recall (+2.6%↑), while still maintaining superior parameter efficiency compared to RFLA. This advancement hinges on auxiliary metadata-guided modulation, which dynamically refines feature representations for robust detection across omni-domain IRSTD challenges.
4.3 Ablation Study
4.3.1 Ablation Study on Multi-Modal Auxiliary Feature
We dissect the contribution of multi-modal feature engineering through three progressive configurations (Tab. 2), revealing fundamental design insights:
1) Cross-Modal Context Priming: Integrating metadata via M2DM elicits 2.4% AP50 and 2.0% Recall gains (Strategy a→b). This stems from metadata’s capacity to resolve infrared imaging ambiguities: spectral bands disambiguate thermal clutter, while sensor parameters guide adaptive receptive field scaling for small targets.
2) Semantic Overload Phenomenon: Directly incorporating deep semantic features causes 0.4% AP50 degradation (Strategy b→c). We attribute this to channel entanglement between high-level semantics (object-centric patterns) and auxiliary metadata (sensor physics), diluting spatial precision that is critical for small targets.
3) Differential Detail Amplification: Embedding compensation detail features in the auxiliary features (Strategy d) further improved AP50 by 0.1% and Recall by 0.2%, highlighting the importance of differential detail information for IRSTD.
Table 2: Ablation Study on Auxiliary Feature Composition
Setting AP50 (%) $\uparrow$ Recall (%) $\uparrow$
Baseline (No Aux.) 74.7 84.5
+ Metadata 77.1 86.5
+ Metadata + High-Level 76.7 86.2
+ Metadata + Detail Comp. 77.2 86.7
4.3.2 Ablation Study on Multi-Modal Dynamic Modulation
Comparative analysis of PSSM/PSCM integration strategies (Tab. 3) exposes the following findings:
1) Independent Enhancement Pathways: PSSM’s edge-sensitive spatial warping (+1.1% AP50) compensates for target boundary erosion in thermal diffusion scenarios. PSCM’s cross-channel contextual modeling increases the performance by 1.0% (74.7% $\rightarrow$ 75.7%).
2) Parallel Fusion Superiority: Parallel fusion (Strategy (f)) achieves 77.2% AP50 via complementary interaction, namely spatial modulation preserves structural integrity while channel attention suppresses clutter harmonics.
3) Parameter-Efficient Design: Shared dynamic projection layers reduce duplicate parameters, achieving 1.84M overhead (only 4.1% of baseline).
Table 3: Ablation Study on Module Combinations and Parameter Analysis
 ID Module Config. Performance Params. (M)
Architecture AP50 (%) $\uparrow$ Recall (%) $\uparrow$
(a) Baseline 74.7 84.5 42.527
(b) PSSM Only 75.8 85.7 44.234
(c) PSCM Only 75.7 85.3 44.230
(d) PSSM $\rightarrow$ PSCM 76.3 86.1 44.367
(e) PSCM $\rightarrow$ PSSM 76.8 86.3 44.367
(f) PSSM + PSCM 77.2 86.7 44.367
4.3.3 Ablation Study on Lightweight Edge Enhancement
To substantiate the efficacy and efficiency of the LEEM module, we benchmarked it against a $3\times 3$ edge convolution operator, initialized to mimic LEEM’s behavior (per Eq. (17)). As reported in Tab. 4, our decomposed 1D convolutional design delivers commensurate performance gains with a parameter increase of less than 1%. This stems from factorizing 2D convolutions into separable 1D operations, which slashes computational overhead while retaining edge-sensitive feature extraction.
Table 4: Ablation study on Lightweight Edge Module vs $3\times 3$ Standard Convolution
Module Structure Performance  Params. (M)
Kernel AP50 (%) $\uparrow$ Recall (%) $\uparrow$
Baseline 74.7 84.5 42.527
W3×3 (Standard) 77.3 86.7 43.116

### v2

@@ III-A Overall Architecture
As illustrated in Fig. 4, the proposed AuxDet framework establishes a novel multimodal architecture for omni-domain IRSTD, building upon the Cascade RPN [37] as the baseline. Our architecture introduces three dedicated components that synergistically address cross-domain generalization challenges: (1) An auxiliary metadata processing stream that transforms heterogeneous sensor parameters into latent semantic representations, (2) A hierarchical fusion mechanism for vision-metadata co-adaptation, and (3) A edge-aware refinement pipeline for target detail recovery.
@@
The pipeline initiates with parallel feature extraction pathways, a conventional CNN backbone for visual pattern mining and a metadata pre-encoder that embeds physical sensor characteristics into high-dimensional manifolds through learnable tensor projections. These complementary representations undergo alignment in Multi-Modal Dynamic Modulation Module. The modulated features then enter our Lightweight Edge Enhancement Module, which deploys phased 1D convolutional blocks along spatial-channel dimensions to amplify high-frequency target signatures.
Figure 5: Architecture of the M2DM module.
III-B Auxiliary Feature Processing and Fusion
@@
Given an input image $\mathbf{x}\in\mathbb{R}^{3\times H\times W}$, we extract hierarchical visual features $\{\mathbf{x}_{i}\}_{i=1}^{4}$ through the backbone, where shallow layers ($i=1,2$) preserve critical spatial details for small target localization, and thus become the embedded layers for our core modules. Our multimodal framework introduces parallel processing of heterogeneous auxiliary metadata $\mathcal{M}=\{platform,band,resolution\}$ via the Multi-Modal Dynamic Modulation module, establishing cross-modal correspondences in the latent space.
III-B1 Heterogeneous Metadata Encoding
@@
To address the multimodal nature of auxiliary information $\mathcal{M}$, we devise a structured encoding strategy. Categorical variables undergo one-hot encoding followed by modality-specific projection:
@@
where $\sigma_{w},\sigma_{h}$ represent dataset-specific normalization factors. The multimodal embeddings undergo modality projection operations to align with visual feature space:
@@
III-B2 Cross-Modal Feature Compensation
We believe that due to the characteristics of small targets, detection models tend to focus on shallow information, with insufficient utilization of deep semantic information. Although auxiliary feature information can provide global semantic information to the model and serve as implicit cues for detection tasks, the edge and detail information of small targets still needs further enhancement. In other words, we need to explore the potential of the multimodal path in compensating for the lack of deep-level detail image features.
Based on this idea, we propose a layer-by-layer differential detail amplifier. As illustrated in Fig. 4, shallow features $\textbf{x}_{i}$ undergo strided convolution $\mathcal{C}_{\downarrow}$ for spatial alignment, generating pseudo-deep features that preserve high-frequency gradients:
@@ III-B2 Cross-Modal Feature Compensation
This design explicitly captures geometric deformation information across layers and aggregates high-frequency detail components from shallow networks with information expression from deep networks.
@@
$\textbf{x}_{e}=\mathcal{F}_{e}(\Delta\textbf{x}),\quad\mathcal{F}_{e}\in\mathbb{R}^{C\times C/4}.$ (6)
@@
where $\mathbf{\Psi}_{aux}$ implements three-stage residual MLP blocks.
This fusion mechanism has dual advantages: the global semantic feature provides contextual priors and decision-level assistance for detection, while the compensation detail feature strengthens the network’s response to subpixel-level edge features through cross-layer gradient propagation, effectively alleviating the dilution of detail information in deep features.
III-C Multi-Modal Dynamic Modulation Module
As shown in Fig. 5, our module establishes an instance-adaptive fusion paradigm through conditional feature transformation. Given aligned visual feature $\mathbf{x}_{i}$ and auxiliary feature $\mathbf{a}_{i}$ , we first project them into latent interaction space via average pooling and learnable linear bottleneck $\mathcal{P}_{\theta}\in\mathbb{R}^{C\times D}$ where $D=\lfloor C/2\rfloor$, then progressively refine the representations through two-stage dynamic MLP transformations [35]. This cascaded architecture enables iterative cross-modal attention:
$\mathbf{y}^{\prime}_{i}=\mathcal{F}_{dyn}^{(2)}\left(\mathcal{F}_{dyn}^{(1)}(\mathcal{P}_{\theta}(\mathbf{y}_{i}))\right),$ (8)
@@ III-C Multi-Modal Dynamic Modulation Module
Unlike conventional batch-shared modulation, our per-sample parameterization preserves batch dimension $B$ to enable instance-wise domain adaptation based on metadata.
@@
$\mathbf{x}_{i}^{c}=\mathbf{x}_{i}\otimes\sigma(\mathbf{cm}_{i}),$ (10)
where $\otimes$ denotes Hadamard product and $\sigma(\cdot)$ represents the sigmoid activation function.
@@
$\mathbf{f}_{sm_{i}}=\sum_{k=1}^{2}\mathcal{G}_{k}\ast\mathbf{f}_{pool_{i}}^{(k)},\quad\mathcal{G}_{k}\in\mathbb{R}^{G\times 1\times 3\times 3},$ (13)
@@
III-D Lightweight Edge Enhancement Module
A pivotal challenge in IRSTD lies in the vulnerability of small target features to being overwhelmed by intricate background clutter. To counter this, we introduce a Lightweight Edge Enhancement Module aimed at refining fused features and restoring subtle target details, while minimizing the introduction of additional parameters.
@@ III-D Lightweight Edge Enhancement Module
where ${W}_{3\times 1}$ and ${W}_{1\times 3}$ denote learnable convolutional kernels, $\alpha\in\mathbb{R}$ acts as a gating parameter. Inspired by Marr-Hildreth edge detection theory, we initialize kernels with discrete Laplacian values to establish stable geometric priors:
@@
where ${I}_{C}\in\mathbb{R}^{C\times C}$ represents the identity matrix that preserves channel-wise independence during convolution kernel initialization and the 1D Laplacian vector is repeated across all $C$ channels.
This initialization strategy simulates the edge sensitivity of local difference operators. When the vertical kernel ${W}_{3\times 1}$ and horizontal kernel ${W}_{1\times 3}$ are cascaded for convolution, the equivalent 3×3 convolution kernel can be represented as the 2D Laplacian-of-Gaussian (LoG) filtering:
@@
The kernel satisfies the second-order differential properties of the discrete Laplacian operator: the center element value (4/16 = 0.25) equals the negative sum of the values of the surrounding eight neighborhood elements ($(-2)\times 4+1\times 4=-4\Rightarrow-4/16=-0.25$). This mathematical property provides the following advantages:
  1. 1.
    Curvature sensitivity: When the gray difference between the central pixel and the surrounding neighborhood changes nonlinearly (e.g., in edge transition areas), the operator generates a large response. This formulation ensures intrinsic curvature sensitivity, yielding robust edge responses from the initial training stages. Consequently, our model achieves superior small target awareness, outperforming most baselines (AP50: 73.0%+) within just 6 epochs.
  2. 2.
    Zero-sum property: In uniform regions, it satisfies $\sum_{i,j}W_{i,j}=0$, suppressing artifacts caused by lighting changes.
  3. 3.
    Isotropy: The rotationally symmetric kernel structure is insensitive to edge direction, ensuring balanced detection of edges at multiple angles.
Through L2-norm scaling (the $\frac{1}{16}$ factor), the initial gradient norm is constrained, effectively avoiding gradient explosion at the beginning of training.
Then, we apply Lightweight Instance Normalization (LW-IN) per sample to mitigate batch normalization’s domain shift issues, ensuring robust performance across diverse imaging scenarios. The output is expressed as:
@@
where $\mu_{\text{space}}$ and $\sigma_{\text{space}}^{2}$ denote the spatial mean and variance computed across each channel, $\gamma\in\mathbb{R}^{C}$ is the learnable scale parameter, and $\epsilon=10^{-5}$ is used for numerical stability. This design preserves feature distribution stability across heterogeneous imaging conditions by eliminating cross-sample statistical coupling.
IV Experiments
IV-A Experimental Settings
1) Datasets: To thoroughly evaluate the effectiveness of the proposed AuxDet framework, we adopt the publicly available WideIRSTD-Full Dataset [18]. It integrates seven prominent public datasets, namely SIRST-V2 [25], IRSTD-1K [16], IRDST [38], NUDT-SIRST [39], NUDT-SIRST-Sea [40], NUDT-MIRSDT [41], and Anti-UAV [42], augmented by an additional dataset from the National University of Defense Technology. It spans diverse imaging platforms, spectral bands, and target types, encompassing both simulated and real-world data with significant variability, which includes 9,000 training images with corresponding metadata. And we further split 7,000 images for training and 2,000 images for validation.
2) Evaluation Metrics: We employ Recall and Average Precision at IoU threshold 0.5 ($AP_{50}$) as primary metrics.
3) Implementation Details: The proposed method is implemented using PyTorch 2.0.1 and MMDetection 3.3.0, with an NVIDIA RTX 3090 acceleration via CUDA 11.7. The maximum training duration is 12 epochs, with the learning rate initialized at 0.005 using a linear warmup for the first 500 iterations. Given the prevalence of large-scale images in the dataset, the batch size is set as 4 and all images are resized to $1024\times 1024$.
For loss functions: Smooth L1 loss is adopted for bbox regression in the ROI Head. IoU loss serves as the refinement loss for two-stage candidate box regression. CrossEntropy loss handles classification in the second stage.
TABLE I: Comparison with Other State-of-the-art methods on WideIRSTD-Full Dataset.
Method Venue Backbone WideIRSTD-Full Dataset
FLOPs $\downarrow$ Params $\downarrow$ AP50 (%) $\uparrow$ Recall (%) $\uparrow$
General Object Detection    $\triangleright$ One-stage    $\triangleright$ Two-stage    $\triangleright$ End-to-End
[…]
@@ IV-A Experimental Settings
Ours - ResNet50 0.215T 45.279M 77.9 87.2
IV-B Comparison with State-of-the-Arts
We conduct rigorous comparisons with 31 representative detectors, including general object detectors (one-stage, two-stage, and end-to-end) and small-target specialized detectors. As shown in Tab. I, our analysis reveals three principal observations:
@@ IV-B Comparison with State-of-the-Arts
1) Generic Detector Inadequacy: Despite architectural advances, generic detectors for natural images exhibit catastrophic failures (avg. AP50<<45%) due to inherent scale biases and the inability to effectively capture small targets. Among these methods, Cascade RPN emerges as the best performer (74.4% AP50) through its multi-stage anchor refinement and adaptive convolution that mitigates feature misalignment. DINO’s contrastive denoising (82.8% Recall) improves small object sensitivity but suffers from transformer’s intrinsic quadratic memory complexity.
@@
2) Limitations of Single-Modal Specialized Models: Domain-specific approaches reveal critical tradeoffs. For example, EFLNet’s adaptive threshold focal loss (73.7% AP50) successfully suppresses background but over-prunes channel dimensions, collapsing on faint targets (68.1% Recall).
@@
3) AuxDet’s Multi-Modal Breakthrough: Our framework achieves SOTA 77.9% AP50 and 87.2% Recall, while still maintaining superior parameter efficiency compared to EFLNet. This advancement hinges on metadata-guided modulation, which dynamically refines feature representations for robust detection across omni-domain IRSTD.
IV-C Visual Analysis
As illustrated in Fig. 6, we compare our AuxDet with the top-4 performing methods. In scenarios where targets are submerged in background clutter (first row), Cascade RPN and SABL exhibit missed detections, while DINO demonstrates significant bounding box regression errors. For dense low-contrast targets (second and fourth rows), RetinaNet and SABL completely fail to localize small targets, whereas Cascade RPN and DINO suffer from both missed detections and false alarms. In wave-texture interference scenarios (third row), SABL misses partial targets, RetinaNet shows severe false positives, and DINO misidentifies waves as targets. In contrast, our method achieves precise target detection with accurate bounding box regression, delivering visually superior results.
Figure 6: Visualization of detection results comparing AuxDet with RetinaNet, Cascade RPN, SABL, DINO.
In the visualization, we use red and blue masks to indicate clearly distinguishable sea areas and visually ambiguous regions with strong cloud interference in maritime detection scenes, respectively. As observed, when the sea and land areas are relatively balanced and the sea is clearly visible to the naked eye (columns 3–4), our method provides a strong prior localization toward the sea regions. When the sea area is minimal (columns 5–6) and small land masses are visually similar to the targets, our method is still able to directly focus on the actual targets, effectively reducing false alarms. In highly entangled scenarios where sea, atmosphere, and land are difficult to distinguish (columns 7–8), our method not only maintains strong prior localization toward the sea, but also successfully highlights target regions partially obscured by clouds and barely discernible to the human eye.
TABLE II: Evaluation of general detectors and small-target specialized detectors with and without AuxDet.
Method Before (Baseline) After (w/ AuxDet)
General Object Detection
Faster R-CNN [53] 31.2 33.5 34.8 (+3.6) 36.6 (+3.1)
Cascade R-CNN [54] 35.8 37.5 37.5 (+1.7) 39.0 (+1.5)
Dynamic R-CNN [57] 35.0 37.2 36.8 (+1.8) 38.8 (+2.0)
Grid R-CNN [55] 32.2 34.8 33.8 (+1.6) 36.0 (+1.2)
Libra R-CNN [56] 30.7 33.8 32.2 (+1.5) 34.8 (+1.0)
SABL [58] 64.1 77.8 65.9 (+1.8) 79.3 (+1.5)
Small Target Specialized
OSCAR [25] 40.2 65.4 47.5 (+7.3) 71.5 (+6.1)
RFLA [65] 71.8 80.6 73.5 (+1.7) 84.1 (+3.5)
NWD [64] 48.5 53.6 61.1 (+12.6) 69.2 (+15.6)
ALCNet [15] 36.3 69.7 42.2 (+5.9) 74.0 (+4.3)
IV-D Analysis of the Integration Flexibility of the Proposed Framework
To demonstrate the generalizability of our approach, we integrate the proposed AuxDet framework as a general-purpose auxiliary module into several representative detection architectures. Specifically, we consider both generic object detectors (e.g., Faster R-CNN, Cascade R-CNN) and small target–oriented methods (e.g., OSCAR, NWD).
As summarized in Tab. II, AuxDet consistently enhances detection performance across all models, yielding improvements in both AP50 and Recall. These gains are observed regardless of the underlying architectural design or training strategies, reflecting the framework’s broad applicability.
Notably, the performance improvements on small target–specialized detectors further indicate that even methods tailored to small-object characteristics can benefit from metadata-guided adaptation. This confirms that AuxDet introduces complementary capabilities that strengthen domain awareness and target discrimination across diverse sensing conditions.
IV-E Comparison with Transformer-based Methods
While our primary comparisons in this paper are conducted against representative CNN-based approaches, Transformer-based models have recently demonstrated promising potential in the field of infrared small target detection (IRSTD), owing to their strong global modeling capabilities and adaptability to complex visual contexts.
To further validate the effectiveness and generalizability of our proposed framework, we conduct two complementary experiments: (1) we integrate AuxDet into Transformer-based backbones to evaluate its compatibility and performance gains in modern vision architectures; (2) we directly compare our method with several state-of-the-art Transformer-based IRSTD methods.
IV-E1 Integration with Transformer Backbones
We replace the CNN backbone with two popular Transformer architectures—PVTv2 [70] and Swin-T [71]—and evaluate the detection performance before and after applying AuxDet. As shown in Tab. III, AuxDet continues to bring consistent improvements across both backbones, achieving +2.2 AP50 and +3.1 Recall on PVTv2-B3, and +2.1 AP50 and +3.2 Recall on Swin-T. These results further demonstrate the general-purpose nature of AuxDet, which can effectively complement Transformer-based visual representations.
TABLE III: Evaluation of AuxDet with transformer backbones.
Backbone Baseline w/ AuxDet